OpenStudy (anonymous):

If you were to solve the system using the addition method, what would you multiply the first equation by in order to make the x’s add out of this system -2x + 3y = 6 4x – 2y = 8 2 -2 4 none of the above

OpenStudy (anonymous):

2 as2(-2x+3y)=-4x+6y so by when adding -4x+6y with4x-2y x's add out of the system

OpenStudy (anonymous):

Solve the following system of equations by graphing. x – y = 7 x + y = 5 What is the x-value of their intersection? 6 -1 12 none of the above

OpenStudy (anonymous):

its 6 draw a graph between x-y=7and x+y=5 itintersects at(6,0) so here x value is 6